Find the infimum and supremum in each of the following sets of real numbers: S = {x| − x2 + 6x − 3 > 0


Let a be the supremum of a set of real numbers and let ε > 0 be any real number. there is at least one x ∈ S such that

a−ε<x≤a

where S is the set with the given supremum


1
Expert's answer
2021-05-13T13:42:52-0400

First Question:

The quadratic equation defined in the set is

"-x^2 +6x - 3 > 0"

"\implies x^2 -6x +3 <0 \\ \implies x^2 -6x +(-3)^2 +3 <0 + (-3)^2 \\ \implies (x-3)^2 < 6"

"\implies x - 3 < \sqrt{6}" or "x - 3 < - \sqrt{6}"

i.e "x < 3+\sqrt{6}" or "x < 3-\sqrt{6}"

This inequality is in this region

"x \in (3- \sqrt{6}, 3+ \sqrt{6})"

Hence the sup S = "3+\sqrt{6}" and inf S "= 3-\sqrt{6}"


Second question:

Suppose such x does not exist, then we have that

"a - \epsilon \geq x \, \forall x \in A"

Hence "a - \epsilon" is an upper bound of A. And by definition of supremum of A, "a \leq a - \epsilon"

This is a contradiction. Hence, there is at least one "x \in A \ni a - \epsilon <x \leq a"